# Smooth Sailing: Lipschitz-Driven Uncertainty Quantification for Spatial Association

### v2

@@ E.3 Computing an Upper Bound on the Bias of Our Estimand with Wasserstein-$1$ Distance
This contradicts the assumption that $C$ is an upper bound on the supremum. Because $C$ was arbitrary, the right hand side of Eq. 7 is infinite, as desired.
@@
We compute the Wasserstein-$1$ distance by linear programming, see discussion in Section B.2. Scalable upper bounds could also be computed by exhibiting a coupling between the measures (for example by solving an entropy regularized optimal transport problem). See [46, Chapters 3 and 4] for details on computation of Wasserstein distances.
E.4 Proof of Lemma 6
Proof of Lemma 6.
We aim to prove that the interval $[-B-\Delta,B+\Delta]$ is the narrowest $1-\alpha$ confidence interval that is valid for all $b\in[-B,B]$ where $\Delta$ is the solution of $\Phi\left(\Delta\right)-\Phi\left(-2B/\tilde{c}-\Delta\right)=1-\alpha$.
@@ E.4 Proof of Lemma 6
$\mathrm{Pr}(X_{b}\in[-B-\Delta,B+\Delta])=\Phi\left(\frac{B+\Delta-b}{\tilde{c}}\right)-\Phi\left(\frac{-B-\Delta-b}{\tilde{c}}\right),$
@@
This ensures $1-\alpha$ coverage over all possible values of $b$ in $[-B,B]$.
@@
$C(b;\Delta)=\Phi\left(\frac{B+\Delta-b}{\tilde{c}}\right)-\Phi\left(\frac{-B-\Delta-b}{\tilde{c}}\right),$
denote the coverage probability of the interval $[-B-\Delta,B+\Delta]$ for $X_{b}\sim\mathcal{N}(b,\tilde{c}^{2})$. In order to ensure the interval is valid for all $b$ coverage, we want to bound below.
@@
The interval $[-B-\Delta,B+\Delta]$ is symmetric about 00, and the Probability Density Function for a Gaussian of mean $b$ is symmetric about $b$. Thus, the coverage probabilities at $b=-B$ and $b=B$ are equal. Consequently, it suffices to consider $b\in[0,B]$.
Moreover, observe that $C(b;\Delta)$ is a strictly decreasing function of $b$ on $[0,B]$ since (i) $\Phi\left(\frac{B+\Delta-b}{\tilde{c}}\right)$ decreases as $b$ increases (because $B+\Delta-b$ decreases and $\Phi(z)$ is monotonic) and (ii) $\Phi\left(\frac{-B-\Delta-b}{\tilde{c}}\right)$ also decreases as $b$ increases (because $-B-\Delta-b$ becomes more negative). Thus, $C(b;\Delta)$ is a strictly decreasing function of $b$ on $[0,B]$. The minimum value of $C(b;\Delta)$ occurs at $b=B$.
@@
$C(B;\Delta)=\Phi\left(\frac{\Delta}{\tilde{c}}\right)-\Phi\left(\frac{-2B-\Delta}{\tilde{c}}\right).$
To ensure that the interval $[-B-\Delta,B+\Delta]$ achieves at least $1-\alpha$ coverage for all $b\in[-B,B]$, we solve:
@@
Narrowest interval. The narrowest interval corresponds to the smallest $\Delta$ that satisfies the Eq. 62. By construction, any smaller $\Delta$ would fail to achieve the required coverage for $b=\pm B$, violating the validity condition. ∎
E.5 Proof of Proposition 8
In this section, we prove Proposition 8. For simplicity of exposition, we prove the result for $\mathcal{S}=[0,1]^{D}$. The result generalizes to any spatial domain which is a compact metric space.
Proof of Proposition 8.
Using Assumption 2 and expanding the quadratic form $(\epsilon_{n}+(f(S_{n})-g(S_{n}))$, we have
@@ E.5 Proof of Proposition 8
Because $f_{N}$ has empirical mean 00 and is continuous, by the intermediate value theorem it takes on the value 00 somewhere on $[0,1]^{D}$. Because $f_{N}$ is $L$-Lipschitz and defined on a set of diameter $\sqrt{D}$, and 00 somewhere inside this set, $\|f_{N}\|_{\infty}\leq L\sqrt{D}$.
Define the set $\overline{\mathcal{F}_{L}}=\mathcal{F}_{L}\cap B_{\infty}(2L\sqrt{D}+2\sigma^{2})$, where $B_{\infty}(r)$ denotes the space of functions uniformly bounded by constant $r$ on $[0,1]^{D}$. By subadditivity of measure, for any $\delta<0$
@@
We first consider the second term in this sum and show it tends to 00. We apply a crude Cauchy-Schwarz bound to the second term in Eq. 65 so that for any $g$,
@@
Because every Lipschitz continuous function is equicontinuous, functions in $\overline{\mathcal{F}_{L}}$ are pointwise bounded by $2L\sqrt{D}+2\sigma^{2}$, and $[0,1]^{D}$ is compact, we may apply Arzela-Ascoli [53, Theorem 7.25] to conclude that $\overline{\mathcal{F}_{L}}$ with the sup norm is sequentially compact. It is therefore compact, as a sequentially compact metric space is compact.
@@
where is a nonnegative function such that $\lim_{\gamma\to 0}\rho(\gamma)=0$. Therefore, for any $\delta<0$, we can find a $\gamma$ such that $\rho(\gamma)\leq-\frac{\delta}{2}$. For this $\gamma$, applying Eq. 74 with $\delta^{\prime}=\frac{\delta}{2}$ allows us to conclude that
@@
This is a uniform law of large number for the class of Lipschitz and bounded functions. More quantitative results are likely possible using empirical process theory, see Wainwright, [67, Chapter 5]. ∎
Proof of Asymptotic Coverage (Corollary 9).
We now prove Corollary 9. We begin by recalling the definition of an asymptotically valid confidence interval.
@@
Corollary 9 follows from Theorem 7 and Proposition 8 by the following lemma, which is a special case of Slutsky’s lemma.
@@
If $\sigma^{2}=0$, then $\hat{\sigma}^{2}_{N}=0$ for all $N$, because the conditional expectation is an $L$-Lipschitz function leading to 00 squared error in the minimization algorithm used to calculate $\hat{\sigma}^{2}_{N}$. Therefore, the resulting confidence interval has coverage $(1-\alpha)$ for all $N$ by Theorem 7.
For $\sigma^{2}>0$ we apply Lemma 14 to conclude that
@@
Convergence in probability implies convergence in distribution, and therefore convergence of quantiles at continuity points. The Gaussian CDF is continuous. Therefore, the quantile computation in Lemma 6 using $\hat{\sigma}^{2}_{N}$ in place of $\sigma^{2}$ produces an asymptotically valid confidence interval in the limit as $N\to\infty$.

### v3

@@ F.3 Computing an Upper Bound on the Bias of Our Estimand with Wasserstein-$1$ Distance
This contradicts the assumption that $C$ is an upper bound on the supremum. Because $C$ was arbitrary, the right hand side of Eq.˜7 is infinite, as desired.
@@
We compute the Wasserstein-$1$ distance by linear programming, see discussion in Section˜B.2. Scalable upper bounds could also be computed by exhibiting a coupling between the measures (for example by solving an entropy regularized optimal transport problem). See [49, Chapters 3 and 4] for details on computation of Wasserstein distances.
F.4 Proof of Lemma 6
Proof of Lemma˜6.
We aim to prove that the interval $B-\Delta,B+\Delta]$ is the narrowest $1-\alpha$ confidence interval that is valid for all $b\in[-B,B]$ where $\Delta$ is the solution of $\Phi\left(\Delta\right)-\Phi\left(-2B/\tilde{c}-\Delta\right)=1-\alpha$.
@@ F.4 Proof of Lemma 6
[…]
@@
This ensures $1-\alpha$ coverage over all possible values of $b$ in [B,B][-B,B].
@@
[…]
denote the coverage probability of the interval $B-\Delta,B+\Delta]$ for $X_{b}\sim\mathcal{N}(b,\tilde{c}^{2})$. In order to ensure the interval is valid for all $b$ coverage, we want to bound below.
@@
The interval $B-\Delta,B+\Delta]$ is symmetric about 0, and the Probability Density Function for a Gaussian of mean $b$ is symmetric about $b$. Thus, the coverage probabilities at b=Bb=-B and $b=B$ are equal. Consequently, it suffices to consider $b\in[0,B]$.
[…]
@@
[…]
To ensure that the interval $B-\Delta,B+\Delta]$ achieves at least $1-\alpha$ coverage for all $b\in[-B,B]$, we solve:
@@
Narrowest interval. The narrowest interval corresponds to the smallest $\Delta$ that satisfies the Eq.˜62. By construction, any smaller $\Delta$ would fail to achieve the required coverage for $b=\pm B$, violating the validity condition. ∎
F.5 Proof of Proposition 8
In this section, we prove Proposition˜8. For simplicity of exposition, we prove the result for $\mathcal{S}=[0,1]^{D}$. The result generalizes to any spatial domain which is a compact metric space.
Proof of Proposition˜8.
Using Assumption˜2 and expanding the quadratic form $(\epsilon_{n}+(f(S_{n})-g(S_{n}))$, we have
@@ F.5 Proof of Proposition 8
Because $f_{N}$ has empirical mean 0 and is continuous, by the intermediate value theorem it takes on the value 0 somewhere on $0,1]^{D}$. Because $f_{N}$ is $L$-Lipschitz and defined on a set of diameter $\sqrt{D}$, and 0 somewhere inside this set, $\|f_{N}\|_{\infty}\leq L\sqrt{D}$.
Define the set $\overline{\mathcal{F}_{L}}=\mathcal{F}_{L}\cap B_{\infty}(2L\sqrt{D}+2\sigma^{2})$, where $B_{\infty}(r)$ denotes the space of functions uniformly bounded by constant $r$ on $0,1]^{D}$. By subadditivity of measure, for any $\delta<0$
@@
We first consider the second term in this sum and show it tends to 0. We apply a crude Cauchy-Schwarz bound to the second term in Eq.˜65 so that for any $g$,
@@
Because every Lipschitz continuous function is equicontinuous, functions in $\overline{\mathcal{F}_{L}}$ are pointwise bounded by $2L\sqrt{D}+2\sigma^{2}$, and $0,1]^{D}$ is compact, we may apply Arzela-Ascoli [58, Theorem 7.25] to conclude that $\overline{\mathcal{F}_{L}}$ with the sup norm is sequentially compact. It is therefore compact, as a sequentially compact metric space is compact.
@@
where is a nonnegative function such that $\lim_{\gamma\to 0}\rho(\gamma)=0$. Therefore, for any $\delta<0$, we can find a $\gamma$ such that $\rho(\gamma)\leq-\frac{\delta}{2}$. For this $\gamma$, applying Eq.˜74 with $\delta^{\prime}=\frac{\delta}{2}$ allows us to conclude that
@@
This is a uniform law of large number for the class of Lipschitz and bounded functions. More quantitative results are likely possible using empirical process theory, see Wainwright, [73, Chapter 5]. ∎
Proof of Asymptotic Coverage (Corollary˜9).
We now prove Corollary˜9. We begin by recalling the definition of an asymptotically valid confidence interval.
@@
Corollary˜9 follows from Theorem˜7 and Proposition˜8 by the following lemma, which is a special case of Slutsky’s lemma.
@@
If $\sigma^{2}=0$, then $\hat{\sigma}^{2}_{N}=0$ for all $N$, because the conditional expectation is an $L$-Lipschitz function leading to 0 squared error in the minimization algorithm used to calculate $\hat{\sigma}^{2}_{N}$. Therefore, the resulting confidence interval has coverage $(1-\alpha)$ for all $N$ by Theorem˜7.
For $\sigma^{2}>0$ we apply Lemma˜14 to conclude that
@@
Convergence in probability implies convergence in distribution, and therefore convergence of quantiles at continuity points. The Gaussian CDF is continuous. Therefore, the quantile computation in Lemma˜6 using $\hat{\sigma}^{2}_{N}$ in place of $\sigma^{2}$ produces an asymptotically valid confidence interval in the limit as $N\to\infty$.